# Coherent Superconductor-Semiconductor Epitaxy for Integrated Quantum Electronics

### v2

@@ Abstract
Introducing superconductivity into group IV elements by doping has long promised a pathway to introduce quantum functionalities into well-established semiconductor technologies. The non-equilibrium hyperdoping of group III atoms into Si or Ge has successfully shown superconductivity can be achieved, however, the origin of superconductivity has been obscured by structural disorder and dopant clustering. Here, we report the epitaxial growth of hyperdoped Ga:Ge films by molecular beam epitaxy with extreme hole concentrations ($n_{\textup{h}}=4.15\times 10^{21}$ cm-3,  17.9% Ga substitution) that yield superconductivity with a critical temperature of $T_{\textup{c}}=3.5$ K and an out-of-plane critical field of 1 T at 270 mK. Synchrotron-based X-ray absorption and scattering methods reveal that Ga dopants are substitutionally incorporated within the Ge lattice, introducing a tetragonal distortion to the crystal unit cell. Our findings, corroborated by first-principles calculations, suggest that the structural order of Ga dopants creates a narrow band for the emergence of superconductivity in Ge, establishing hyperdoped Ga:Ge as a low-disorder, epitaxial superconductor-semiconductor platform.
preprint: APS/123-QED
@@
Although the realization of superconductivity in group IV elements has been previously demonstrated [1, 2, 3, 4], difficulties such as dopant segregation and precipitation, nanocrystal formation, and incoherent interfaces impede their integration into a new generation of quantum devices. As a prototypical semiconductor and foundry-ready material, germanium is a key component of devices that combine macroscopic superconducting coherence with the microscopic degrees of freedom in semiconductors [5, 6, 7, 8]. While epitaxial superconductor-semiconductor (S-Sm) interfaces have been demonstrated with III-V semiconductors [9, 10], realizing full S-Sm-S heterostructures remains challenging and has limited their applicability in quantum devices [11, 12]. This difficulty has renewed interest in doping-based superconductivity in group IV materials, a concept explored theoretically by Cohen in the 1960s [13], as recent efforts to induce superconductivity through proximity effects [7, 8] and alloying [14] have met with mixed results. This route sidesteps disordered interfaces between dissimilar materials by inducing superconductivity within an epitaxial semiconductor thin film [12], offering a unique and promising solution to the problems of material disorder and device scale-up.
@@
Hyperdoping group IV semiconductors induces a superconducting state, yet routes to high quality, device-ready films have remained out of reach. Multiple studies have confirmed the existence of superconductivity [15, 2, 3, 16], however, these materials suffer from structural inhomogeneities, dopant clustering, and poor thickness control, which limit their applications. In terms of growth, the requirement of dopant concentrations above the standard miscibility limits raises concerns over the structural stability, site disorder, and nature of the superconducting phase [3]. Recent efforts in Ge have pushed dopant concentrations above miscibility limits, but maintaining coherent epitaxy and low defect densities remains challenging as most doping strategies rely on high-energy techniques such as ion implantation and flash annealing. These approaches demonstrate superconductivity but the formation of lattice defects and broad dopant profiles complicate reproducible film growth. This leads to persistent questions about whether these hyperdoped films are intrinsically superconducting, or if Ga clusters and unintended phases are responsible for superconductivity. These uncertainties highlight the need for growth strategies that introduce robust superconductivity into semiconductors without sacrificing crystallinity.
Figure 1: Superconductivity in germanium by p-type hyperdoping. A Schematic picture of dopants within a matrix and zone-center E(k) dispersion for a prototypical band insulator in a dilute doping and hyperdoped state. In the dilute doping scenario, dopants with an effective Bohr radius, $a_{0}$, at an average distance between dopants, $\lambda_{d}$, induce mid-gap electronic states at an energy $E_{a}$ above the valence band edge. Increasing dopant concentrations to few-percent levels or higher induce significant hybridization between dopants, giving rise to a low-energy band associated with dopant-dopant hybridization. This defect band is traditionally expected to be non-dispersive due to the stochastic behavior of dopants. However, dopant-dopant hybridization in principle can induce ordering of the dopants on the lattice, giving rise to a narrow heavy band as well. B and C present sheet resistance versus temperature traces for a single Ga:Ge layer and a Ga:Ge sandwich structure, akin to a Josephson junction along the growth direction, with a thin Si spacer layer, respectively. In both of these structures, the Ga:Ge hyperdoped layer is kept nominally constant, with the single layer thickness at roughly 12 nm and the Ga:Ge layer thicknesses in the sandwich are both 10 nm.
@@
Here, we use molecular beam epitaxy (MBE) to grow Ga-hyperdoped Ge films, overcoming challenges associated with high dopant concentrations while maintaining coherent epitaxial growth. Crucially, the surfactant behavior of Ga enables the low-temperature growth of Ge, preserving smooth surfaces and single-crystallinity [17]. For films with an expected Ga doping concentration of 17.9%, we observe a superconducting transition at 3.5 K and a carrier concentration of 4.15 $\times$ 1021 holes cm-3. Employing short- and long-range structural characterization methods, we show that Ga substitution is energetically favored within the Ge lattice, slightly expanding its unit cell and producing a tetragonal distortion reminiscent of the $\beta$-Sn phase. Density functional theory (DFT) calculations reveal a pronounced shift of the Fermi level into the valence band and a flattening of electronic bands near the R-point of the cubic Brillouin zone. Our findings establish that superconductivity originates intrinsically in Ge from substitutional Ga doping and flat- or narrow-band physics, rather than defect-mediated superconducting weak links.
@@
Figure 1A presents a schematic diagram of the dopant structure under dilute doping conditions, as well as a generalized picture for a hyperdoping in a semiconductor, along with their respective schematic E(k) dispersions. For dilute doping, the average distance between Ga dopants, $\lambda_{d}$, is larger than the effective Bohr radius of the Ga dopant, $a_{0}$. This results in discrete mid-gap states associated with these dopants at an energy $E_{\textup{a}}$ above the valence band edge, as is expected for a hole-doped semiconductor. However, the scenario in which dopant concentrations greatly exceed common doping levels, known as “hyperdoping”, leads to significant dopant-dopant and dopant-semiconductor hybridization of the low energy states. This gives rise to what is typically considered to be a non-dispersive band of defect states. Alternatively, hybridization with the host matrix can indeed give rise to a new heavy electronic band that disperses in $k$. Transport data for MBE-grown samples for this study is shown in Figure 1B and C, for structures containing a single layer of hyperdoped Ge and an epitaxial Josephson junction-like structure, respectively. Compared to previous reports on MBE-grown samples [16], these samples exhibit larger carrier concentrations that are more comparable to ion implantation strategies [15, 2, 3] (See Figure S1), albeit with a slight reduction in superconducting $T_{\textup{c}}$. We note a rough trend in $T_{\textup{c}}$ as a function of hole carriers where superconductivity appears near $1\times 10^{21}$ cm-3, reaching a maximum near $4-5\times 10^{21}$ cm-3. This is consistent with the idea that increasing the density of states (DOS) at the Fermi level increases the superconducting $T_{\textup{c}}$. Furthermore Figure 1B and C present the sheet resistance as a function of temperature for a 12 nm thick film of hyperdoped Ge and a sandwich heterostructure comprised of 10 nm layers of hyperdoped Ge surrounding a thin Si interlayer. The magnetic field dependence of longitudinal resistance within the Josephson structure is presented in Figure S3A and B for out-of-plane and in-plane field configurations, respectively. At 300 mK we observe the out-of-plane $H_{\textup{c}}$ to be 0.79 T, decreasing to 0.14 T at 2 K. In-plane $H_{\textup{c}}$ is found to be 2.51 T, decreasing to 0.78 T at 2 K. The robustness of this superconducting state in a magnetic field makes this superconducting system highly favorable for experiments investigating Andreev bound state behavior at values of magnetic field relevant for topological superconductivity [18, 19].
.1 Dopant fine structure and band structure
To examine the role of Ga as a dopant, we first study the thermodynamic preference and site-specific effects of introducing Ga into the diamond-type Ge crystal using first-principles calculations based on DFT, as detailed in the Methods section. Approximating from our Hall measurements and MBE growth conditions, doping concentrations of 12.5% and 11.1% are simulated for the case of substitutional and interstitial doping, respectively. Figure 2A shows the DFT-optimized structures. Incorporating group-III Ga as a heterovalent dopant within Ge implies some degree of non-equilibrium growth (i.e., the MBE growth process). This results in positive defect formation energies for cells containing Ga located at both doping sites (Figure 2A and Table S1). However, the formation energy of the substitutional Ga-doped structure (0.314 eV) is markedly lower than the interstitial positions (1.946 eV), indicating that substitutional dopants are expected to be more energetically favorable in the material. This preference is inherently connected to the requirement of the Ge host to undergo an expansion to accommodate the larger Ga atom at the interstitial site, as reported in Table S1. From our DFT calculations, substituting Ge with Ga at these doping levels is expected to impart a relatively small out-of-plane expansion of the lattice parameter of $\sim$0.12%, resulting in a slight tetragonal distortion to the structure (Table S1) to accommodate the heterovalent group-III dopant in a lattice-matched film.
@@
Figure 2: Fine structure measurements revealing dopant locations and modulated band structure. A Optimized DFT structures of Ga-doped Ge at the interstitial and substitutional sites, along with the defect formation energies ($E_{\textup{Form}}$). B Experimental Ga-K-edge spectrum recorded from the Ga:Ge with its corresponding $k^{2}$-weighted $\chi$(k) EXAFS fit (red line) contained in the inset. C The Fourier-transform (k2-weighted) of the EXAFS data (distances have been phase-corrected). An agreeable fit has been made to the experiment using the substitutional doped crystallographic information file (CIF) generated from the optimized DFT calculated structure shown in A.D Calculated band structure for pristine Ge and the substitutionally doped GaGe7 structure. The horizontal dashed red line indicates the position of the Fermi level after doping, shifting 1.01 eV into the valence band. E The contribution of each band identified as i, ii, and iii to the 3D Fermi surface with the color map corresponding to the Fermi velocity. The predicted narrow-band condition is observed as small pockets of high-mass carriers in the surface of Band iii at the corners of the Brillouin Zone (R point).
@@
To verify our theoretical calculations of doping site preference in hyperdoped Ge, we perform synchrotron X-ray absorption spectroscopy measurements. From our calculations (Figure S12), X-ray absorption at the Ga K-edge is inherently sensitive to the Ga-related local fine structure and can be used to unambiguously resolve its location within the lattice. Experimentally, we employ extended X-ray absorption fine structure (EXAFS) analysis, as measured from the Ga:Ge thin film in a fluorescent mode. Multiple spectra were recorded across the sample, with their average normalized spectrum forming the X-ray absorption data shown in Figure 2B. The $k^{2}$-weighted EXAFS signal is shown in the inset, with the corresponding analysis of the Fourier transform presented in Figure 2C. Considering the different structural environments that Ga can occupy in Ge (Figure 2A), the EXAFS data are best described by a substitutional doping model with a strong Ge-Ga nearest-neighbor scattering pair at 2.447 Å, which is close to the calculated bond length of 2.439 Å (Figure 2A). To rule out other possible coordination environments, such as $\alpha$-Ga segregation and the interstitial lattice position, Figure S12 contains the simulated profiles of these fine-structure environments. Together, with no strong fine structure observed between 2.5 and 3.8 Å in the experimental Fourier transform, these findings confirm that the Ga dopants preferentially occupy substitutional lattice positions, in agreement with our ab initio calculations. Furthermore, the relatively large Debye–Waller factors ($\sigma^{2}$( Å2)) derived from the EXAFS fits (Table LABEL:exafs-table) highlight the role of both lattice strain and Si diffusion in the Ga:Ge layers.
@@
With confirmation of substitutional doping, we use the optimized structures in Figure 2A to study how the electronic structure is perturbed through relatively large incorporations of heterovalent Ga dopants at the different lattice sites. While a slight tetragonal distortion is anticipated in strained material due to substrate clamping, we find that the calculated band structure is insensitive to this distortion and we use the high-symmetry cubic Brillouin zone and paths for simplicity. The $k$-resolved band structure is presented in Figure 2D and the $k$-integrated DOS is in Figure S7A. From our formation energy calculations, Ga dopants are energetically unlikely to reside at interstitial lattice positions. This is confirmed by the DOS calculation shown in Figure S7A, which indicates that interstitial doping is undesirable for preserving the “clean” electronic bands of the host Ge semiconductor. This is evidenced by the disappearance of the parent Ge band gap and the emergence of metallic states. In contrast, substitutional Ga doping is expected to form acceptor states below the valence band edge, while the overall band structure near the gap remains largely unperturbed (Figure 2D). An evaluation of the DOS projected onto different atomic orbitals in Figure S7B shows that new, localized states emerge through the hybridization of both the Ga and Ge p-orbitals. As expected, substituting Ge for group III Ga induces p-type doping in the system, with the Fermi level shifting 1.01 eV into the valence band (see Figure 2D). This corresponds to a hole carrier concentration of $5.5\times 10^{21}$ cm-3 that closely approximates the hole concentration from our Hall measurements. Furthermore, due to the lifting and flattening of the bands near the R point of the cubic Brillouin zone, we predict this system to be a candidate narrow-band superconductor. By plotting the Fermi surface and separating the contribution of each of the three bands crossing the Fermi level (band i, ii, and iii), we observe in the surface for band iii, hole pockets of high-mass carriers that cross the Fermi level at the corners of the Brillouin zone (Figure 2E). This feature is clearly represented in the 2D projected cross-sections of the doped Ga:Ge Fermi surface (Figure S8) and is highly sensitive to the position of the Fermi level. Lowering it by 40 meV to simulate a slight increase in the hole concentration will both enlarge the corner hole pocket projections for band iii and introduce similar features into the projected corners of band ii (Figure S9).
@@
Figure 3: Cross-sectional electron microscopy reveals coherent crystalline interfaces. A Schematic of the film structure as grown by MBE. B-C Cross-sectional TEM/EDS imaging of the $T_{\textup{c}}$ = 3.5 K hyperdoped sample presenting the compositional profile across the entire film thickness. D Low-magnification cross-sectional HAADF-STEM image of the film. E Zoom-in on the film/substrate and Ga:Ge/Si interfaces, displaying coherent crystalline interfaces. F Raman back-scattering spectra of undoped (black) and doped (red) structures using 532 nm laser excitation at room temperature. The inset highlights the common emergence of highly anharmonic second-order optical bands, identified as the active LO mode here due to the Raman selection rules for backscattering from a (001) Ge surface.
.2 Epitaxial interfaces and Phonons
@@
Figure 3A presents a schematic structure of the grown film with $T_{\textup{c}}$ = 3.5 K, in which two hyperdoped Ge layers are grown sandwiched between a thin Si buffer layer. Cross-sectional transmission electron microscopy (TEM) and energy dispersive spectroscopy (EDS) measurements are presented in Figure 3B-C and show the spatially homogeneous elemental profile of the Ga dopants throughout the hyperdoped layer. Notably, no distinct signatures of Ga clustering are observed compared to earlier reports where annealing was used to activate dopants  [3]. Likewise, these MBE-grown materials also exhibit abrupt interfaces and a smooth surface amenable to future planar device fabrication as opposed to previous MBE studies [16]. A view of the atomic structure using low-magnification high-angle annular dark field scanning transmission electron microscopy (HAADF-STEM) is presented in Figure 3D, with a corresponding expansion of the key film/substrate and Ga:Ge/Si interfaces contained in Figure 3E. Owing to the small amount of mismatch strain and thin layer thicknesses, the growth of Ga:Ge is pseudomorphic (commensurate) with the underlying Ge substrate and the atomic planes on both sides of the interface are in perfect registry. This demonstration of fine thickness control, commensurate interfaces, and homogeneous dopant distribution exhibits all of the necessary structural requirements for the fabrication of low-disorder superconducting devices that have been proposed to mitigate loss channels in quantum information applications [12].

### v3

# Superconductivity in Substitutional Ga-Hyperdoped Ge Epitaxial Thin Films

@@ Abstract
Doping-induced superconductivity in group IV elements may enable quantum functionalities in material systems accessible with well-established semiconductor technologies. Non-equilibrium hyperdoping of group III atoms into C, Si, or Ge can yield superconductivity; however, its origin is obscured by structural disorder and dopant clustering. Here, we report the epitaxial growth of hyperdoped Ga:Ge films and trilayer heterostructures by molecular beam epitaxy with extreme hole concentrations ($n_{\textup{h}}=4.15\times 10^{21}$ cm-3,  17.9% Ga substitution) that yield superconductivity with a critical temperature of $T_{\textup{c}}=3.5$ K. Synchrotron-based X-ray absorption and scattering methods reveal that Ga dopants are substitutionally incorporated within the Ge lattice, introducing a tetragonal distortion to the crystal unit cell. Our findings, corroborated by first-principles calculations, suggest that the structural order of Ga dopants creates a narrow band for the emergence of superconductivity in Ge, establishing hyperdoped Ga:Ge as a low-disorder, epitaxial superconductor-semiconductor platform.
@@
Realization of superconductivity in group IV elements has been an enticing field of study over the past few decades for their promising application in superconducting electronics Ekimov et al. (2004); Bustarret et al. (2006); Blase et al. (2009); Chiodi et al. (2017) such as superconducting quantum bits or cryogenic CMOS control circuitry. However, difficulties such as dopant segregation and precipitation, nanocrystal formation, incoherent interfaces, and poor layer thickness control impede their integration into a new generation of quantum devices. As a prototypical semiconductor and foundry-ready material, germanium is a key component of devices that combine macroscopic superconducting coherence with the microscopic degrees of freedom in semiconductors Scappucci et al. (2021); Sammak et al. (2019); Tosato et al. (2023); Schiela et al. (2024). Historically, superconductivity in germanium has been accomplished through extreme non-equilibrium growth techniques such as ion implantation and subsequent flash annealing Hermannsdörfer et al. (2009); Prucnal et al. (2019); Sardashti et al. (2021). However, recent advancements in thin film epitaxy of heavily doped germanium thin films have shown great promise in achieving well-ordered, single-crystalline thin films of superconducting germanium Strohbeen et al. (2023). Homoepitaxy of superconducting group IV materials allows for expanding upon the existing repertoire of superconductor-semiconductor (S-Sm) planar devices Shabani et al. (2016); Krogstrup et al. (2015) to enable full S-Sm-S heterostructures that have remained challenging to realize, thus limiting their applicability in quantum devices Zhao et al. (2020); Shim and Tahan (2014). This difficulty has renewed interest in doping-based superconductivity in group IV materials, a concept explored theoretically by Cohen in the 1960s Cohen (1964), as recent efforts to induce superconductivity through proximity effects Tosato et al. (2023); Schiela et al. (2024) and alloying Strohbeen et al. (2024) have met with mixed results. This route sidesteps disordered interfaces between dissimilar materials by inducing superconductivity within an epitaxial semiconductor thin film Shim and Tahan (2014), offering a unique and promising solution to the problems of material disorder and device scale-up.
@@
Hyperdoping in germanium induces a superconducting state, yet routes to high quality, device-ready films have remained out of reach. Multiple studies have confirmed the existence of superconductivity Hermannsdörfer et al. (2009); Prucnal et al. (2019); Sardashti et al. (2021); Strohbeen et al. (2023), however, these materials suffer from structural inhomogeneities, dopant clustering, and poor thickness control, which limit their applications. Specifically, doping at concentrations above the standard miscibility limits raises concerns over the structural stability, site disorder, and nature of the superconducting phase Sardashti et al. (2021). Recent efforts in Ge have pushed dopant concentrations above miscibility limits Strohbeen et al. (2023); Trumbore (1960); Sardashti et al. (2021); Prucnal et al. (2019); Hermannsdörfer et al. (2009), but maintaining coherent epitaxy and low defect densities remains challenging as most doping strategies rely on high-energy techniques such as ion implantation and flash annealing. These approaches demonstrate superconductivity but the formation of lattice defects and broad dopant profiles complicate device integration and dopant uniformity. Furthermore, questions persist about whether these hyperdoped films are intrinsically superconducting, or if Ga clusters and unintended phases are responsible for superconductivity. These uncertainties highlight the need for growth strategies more readily amenable to achieve the quality necessary for quantum devices as well as better understanding of dopant incorporation in these extreme doping regimes.
@@
Here, we use molecular beam epitaxy (MBE) to grow Ga-hyperdoped Ge films, overcoming challenges associated with high dopant concentrations while maintaining coherent epitaxial growth. Crucially, better control over parasitic heating during film growth enables the formation of a smooth floating layer of Ga adatoms that promotes low-temperature growth of Ge, preserving smooth surfaces and single-crystallinity Kesan et al. (1991). By re-configuring the growth chamber and optimizing growth temperatures, we obtain drastic improvements in surface morphology, atomic ordering, and domain structure. We demonstrate the improved material quality in the growth of trilayer heterostructures as a proof-of-principle design for vertical Josephson junction device architectures that alleviate concerns of material disorder related to typical amorphous oxide tunnel junctions. For films with an expected Ga doping concentration of 17.9%, we observe a superconducting transition at 3.5 K and a carrier concentration of $4.15\times 10^{21}$ holes cm-3. Short- and long-range structural characterization methods show that Ga substitution is energetically favored within the Ge lattice and drives a tetragonal distortion reminiscent of the $\beta$-Sn phase. Density functional theory (DFT) calculations reveal a pronounced shift of the Fermi level into the valence band and a flattening of electronic bands near the R-point of the cubic Brillouin zone. Our findings establish that superconductivity originates intrinsically in Ge from substitutional Ga doping and flat- or narrow-band physics, rather than defect-mediated superconducting weak links.
Figure 1: Superconductivity in germanium by p-type hyperdoping. a Schematic picture of dopants within a matrix and zone-center $E(k)$ dispersion for a prototypical band insulator in a dilute doping and hyperdoped state. In the dilute doping scenario, dopants with an effective Bohr radius, $a_{0}$, at an average distance between dopants, $\lambda_{d}$, induce mid-gap electronic states at an energy $E_{a}$ above the valence band edge. Increasing dopant concentrations to few-percent levels or higher induces significant hybridization between dopants, giving rise to a low-energy band associated with dopant-dopant hybridization. This defect band is traditionally expected to be non-dispersive due to the stochastic behavior of dopants. However, dopant-dopant hybridization in principle can induce ordering of the dopants on the lattice, giving rise to a narrow heavy band as well. b-e Sheet resistance versus temperature traces for single Ga:Ge layers and Ga:Ge sandwich structures, akin to a Josephson junction along the growth direction. In these structures, the single Ga:Ge layers are held to a constant thickness of roughly 12 nm and the trilayer structures have Ga:Ge layer thicknesses of 10 nm. The insets of each panel present schematic drawings of the sample layer structure. b, c are representative traces where we use Si as the cap and barrier layer material, and d, e are representative traces where Ge replaces the Si-containing layers. The jitter in temperature observed in d during the superconducting transition is an artifact caused by temperature instability during the measurement.
Thin film growth and material stability
@@
Figure 1a presents a schematic diagram of the dopant structure under dilute doping conditions, as well as a generalized picture for hyperdoping in a semiconductor, along with their respective schematic $E(k)$ dispersions. For dilute doping, the average distance between Ga dopants, $\lambda_{d}$, is larger than the effective Bohr radius of the Ga dopant, $a_{0}$. This results in discrete mid-gap states associated with these dopants at an energy $E_{\textup{a}}$ above the valence band edge, as is expected for a hole-doped semiconductor. However, the scenario in which dopant concentrations greatly exceed common doping levels, known as “hyperdoping”, leads to significant dopant-dopant and dopant-semiconductor hybridization of the low-energy states. This gives rise to what is typically considered to be a non-dispersive band of defect states. Alternatively, hybridization with the host matrix can indeed give rise to a new heavy electronic band that disperses in $k$. Transport data for MBE-grown samples for this study are shown in Figure 1b and c, for structures containing a single layer of hyperdoped Ge and an epitaxial Josephson junction-like structure, respectively. Compared to previous reports on MBE-grown samples Strohbeen et al. (2023), these films exhibit a slight reduction in superconducting $T_{\textup{c}}$, but larger carrier concentrations that are more comparable to ion implantation strategies Hermannsdörfer et al. (2009); Prucnal et al. (2019); Sardashti et al. (2021) (see Figure S1a). Figure S1a shows a rough trend in $T_{\textup{c}}$ as a function of hole carriers where superconductivity appears near $1\times 10^{21}$ cm-3, reaching a maximum near $4-5\times 10^{21}$ cm-3 in the samples grown for this study. This is consistent with a previous report in boron-doped silicon Nath et al. (2024) that demonstrates an increasing hole density increases the superconducting $T_{\textup{c}}$. Furthermore, Figure 1b and c present the sheet resistance as a function of temperature for a 12 nm thick film of hyperdoped Ge and a trilayer heterostructure, respectively. Representative reflection high-energy electron diffraction (RHEED) images for a layered heterostructure of hyperdoped Ge films are presented in Figure S2. During Ga doping, we observe the formation of the expected floating surface layer of Ga adatoms Kesan et al. (1991) with an underlying single crystalline pattern that reflects the surface structure of the hyperdoped film. Further details regarding RHEED monitoring and film growth are presented in the Supplementary information and Methods, respectively. The smoothness of the resultant film surface is confirmed with ex-situ atomic force microscopy (AFM) imaging, seen in Figure S1b.
To investigate the role of the silicon interlayer more explicitly, films grown without silicon were also studied. The results of which are presented in Figures 1d, e, and S3. We note that the silicon-containing samples stabilize slightly higher carrier concentrations, which we expect to exist due to the band offset between the hyperdoped Ge and strained Si layers Mönch (1987); Van de Walle (1988). Silicon incorporation also enhances the resultant superconducting $T_{\textup{c}}$, which we attribute to a behavior akin to the BCS isotope effect Blase et al. (2009); Nath et al. (2024). Temperature stability of the superconducting phase is shown in Figure S3, where we obtain a maximal superconducting transition temperature at growth temperatures ranging from 100-150C. Above 150C the superconducting phase weakens as observed by a reduction in $T_{\textup{c}}$, coinciding with a reduction in hole carriers seen in Figure S4. We further investigate the robustness of the superconducting phase by fabricating a Hall bar device out of the sample shown in Figure 1c. We present the data for this test device in Figure S5 and the fabrication procedure is found in the Methods. We note no significant change in superconducting $T_{\textup{c}}$, out-of-plane $H_{\textup{c}}$, nor carrier concentration. An extended discussion of this data is provided in the Supplementary Information. Thus, we conclude that this superconducting phase is demonstrably stable during processing and suitable for device applications.
@@
Figure 2: Ga dopant arrangement in Ge and modulated band structure. a Optimized DFT structures of Ga-doped Ge at the interstitial and substitutional sites, along with the defect formation energies ($E_{\textup{Form}}$). b Experimental Ga-K-edge spectrum recorded from the Ga:Ge with its corresponding $k^{2}$-weighted $\chi(k)$ EXAFS fit (red line) contained in the inset. c The Fourier-transform ($k^{2}$-weighted) of the EXAFS data (distances have been phase-corrected). An agreeable fit has been made to the experiment using the substitutional doped crystallographic information file (CIF) generated from the optimized DFT calculated structure shown in a.d Calculated band structure for pristine Ge and the substitutionally doped GaGe7 structure. The horizontal dashed red line indicates the position of the Fermi level after doping, shifting 1.01 eV into the valence band. e The contribution of each band identified as i, ii, and iii to the 3D Fermi surface with the color map corresponding to the Fermi velocity. The predicted narrow-band condition is observed as small pockets of high-mass carriers in the surface of Band iii at the corners of the Brillouin Zone (R point).
Dopant arrangement and band structure
Approximating from our Hall measurements and MBE growth conditions, doping concentrations of 12.5% and 11.1% are simulated for the case of substitutional and interstitial doping, respectively, as detailed in the Methods section. Figure 2a shows the DFT-optimized structures. We find positive defect formation energies for cells containing Ga located at both doping sites (Figure 2a and Table S1), implying some degree of non-equilibrium growth necessary (i.e., MBE growth) in both cases. However, the formation energy of the substitutional Ga-doped structure (0.314 eV) is markedly lower than the interstitial positions (1.946 eV), indicating that substitutional dopants are expected to be more energetically favorable in the material. This preference is inherently connected to the requirement of the Ge host to undergo an expansion to accommodate the larger Ga atom at the interstitial site, as reported in Table S1. From our DFT calculations, substituting Ge with Ga at these doping levels is expected to impart a relatively small out-of-plane expansion of the lattice parameter of $\sim$0.12%, resulting in a slight tetragonal distortion to the structure (Table S1) to accommodate the heterovalent group-III dopant in a lattice-matched film.
@@
To verify our theoretical calculations of doping site preference in hyperdoped Ge, we perform synchrotron X-ray absorption spectroscopy measurements which have previously demonstrated strong capabilities for determining the lattice position of dopant species Zhang et al. (2024). From our calculations (Figure S6), X-ray absorption at the Ga K-edge is inherently sensitive to the Ga-related local fine structure and can be used to unambiguously resolve its location within the lattice. Experimentally, we employ extended X-ray absorption fine structure (EXAFS) analysis, as measured from the Ga:Ge thin film in a fluorescent mode. X-ray absorption data is shown in Figure 2b measured from various points across the sample, normalized and then averaged together with the $k^{2}$-weighted EXAFS signal shown in the inset. The corresponding analysis of the Fourier transform of this data is presented in Figure 2c. The EXAFS data are best described by a substitutional doping model the strongest Ge-Ga nearest-neighbor scattering pair at 2.447 Å, in excellent agreement with the calculated bond length of 2.439 Å (Figure 2a). To rule out other viable dopant coordination environments, such as $\alpha$-Ga segregation and the interstitial lattice position, we simulate the expected profile for these alternative fine-structure environments, shown in Figure S7. These simulations, in combination with no significant fine structure observed between 2.5 and 3.8 Å in the experimental Fourier transform, we confirm that the Ga dopants preferentially occupy substitutional lattice positions, in agreement with our ab initio calculations. Furthermore, the relatively large Debye–Waller factors ($\sigma^{2}$( Å2)) derived from the EXAFS fits (Table S2) highlight the role of both lattice strain and Si diffusion in the Ga:Ge layers.
@@
Using the optimized structures in Figure 2a with the known dopant arrangement, we now study how the electronic structure is perturbed through large incorporations of heterovalent Ga dopants. While a slight tetragonal distortion is anticipated in strained material due to substrate clamping, we find that the calculated band structure is insensitive to this distortion and we use the high-symmetry cubic Brillouin zone and paths for simplicity. The $k$-resolved band structure is presented in Figure 2d and the $k$-integrated DOS is in Figure S8a. An evaluation of the DOS projected onto different atomic orbitals in Figure S8b shows that new, localized states emerge through the hybridization of both the Ga and Ge p-orbitals. As expected, Ga substitution induces p-type doping in the system, with the Fermi level shifting 1.01 eV into the valence band (see Figure 2d). This corresponds to a hole carrier concentration of $5.5\times 10^{21}$ cm-3 that closely approximates the hole concentration from our Hall measurements. Furthermore, due to the lifting and flattening of the bands near the R point of the cubic Brillouin zone, we predict this system to be a candidate narrow-band superconductor. By plotting the Fermi surface and separating the contribution of each of the three bands crossing the Fermi level (band i, ii, and iii), we observe in the surface for band iii, hole pockets of high-mass carriers that cross the Fermi level at the corners of the Brillouin zone (Figure 2e). This feature is clearly represented in the 2D projected cross-sections of the doped Ga:Ge Fermi surface (Figure S9) and is highly sensitive to the position of the Fermi level. Lowering it by 40 meV to simulate a slight increase in the hole concentration will both enlarge the corner hole pocket projections for band iii and introduce similar features into the projected corners of band ii (Figure S10).
Epitaxial interfaces and phonon structure
@@
Figure 3: Cross-sectional electron microscopy reveals coherent crystalline interfaces. a Schematic of the film structure as grown by MBE. b-c Cross-sectional TEM/EDS imaging of the $T_{\textup{c}}$ = 3.5 K hyperdoped sample presenting the compositional profile across the entire film thickness. d Low-magnification cross-sectional HAADF-STEM image of the film. e Zoom-in on the film/substrate and Ga:Ge/Si interfaces, displaying coherent crystalline interfaces. f Raman back-scattering spectra of undoped (black) and doped (red) structures using 532 nm laser excitation at room temperature. The inset highlights the common emergence of highly anharmonic second-order optical bands, identified as the active LO mode here due to the Raman selection rules for backscattering from a (001) Ge surface. For more details regarding the Raman spectra, please see Figure S12.
@@
Figure 3a presents a schematic structure of the grown film with $T_{\textup{c}}$ = 3.5 K, in which two hyperdoped Ge layers are grown surrounding a thin Si buffer layer. Cross-sectional transmission electron microscopy (TEM) and energy dispersive spectroscopy (EDS) measurements are presented in Figure 3b-c and show the spatially homogeneous elemental profile of the Ga dopants throughout the hyperdoped layer. Notably, no distinct signatures of Ga clustering are observed compared to earlier reports where annealing was used to activate dopants Sardashti et al. (2021). Likewise, these MBE-grown materials also exhibit abrupt interfaces and a smooth surface amenable to future planar device fabrication as opposed to previous MBE studies Strohbeen et al. (2023). A view of the atomic structure using low-magnification high-angle annular dark field scanning transmission electron microscopy (HAADF-STEM) is presented in Figure 3d, with a corresponding expansion of the key film/substrate and Ga:Ge/Si interfaces contained in Figure 3e. To verify domain structure, a wider view of the STEM lamella is presented in Figure S11 where we see no distinct signs of polycrystalline domain formation nor significant surface roughness. Owing to the small amount of mismatch strain and thin layer thicknesses, the growth of Ga:Ge is commensurate with the underlying Ge substrate and the atomic planes on both sides of the interface are in perfect registry. We note that in the current study, we measure a barrier thickness of $\sim$1 nm whereas we estimate a thickness of 5 nm of Si or 10 nm of Ge to be necessary for the demonstration of Josephson tunneling. However, this demonstration of fine thickness control, commensurate interfaces, and homogeneous dopant distribution exhibits the necessary structural and chemical requirements to reduce the disorder in superconductive quantum devices and thus positions these hyperdoped trilayer structures to better mitigate microwave photon loss in quantum information applications Shim and Tahan (2014).